### v1

# CricaVPR: Cross-image Correlation-aware Representation Learning for Visual Place Recognition

Abstract
Over the past decade, most methods in visual place recognition (VPR) have used neural networks to produce feature representations. These networks typically produce a global representation of a place image using only this image itself and neglect the cross-image variations (e.g. viewpoint and illumination), which limits their robustness in challenging scenes. In this paper, we propose a robust global representation method with cross-image correlation awareness for VPR, named CricaVPR. Our method uses the self-attention mechanism to correlate multiple images within a batch. These images can be taken in the same place with different conditions or viewpoints, or even captured from different places. Therefore, our method can utilize the cross-image variations as a cue to guide the representation learning, which ensures more robust features are produced. To further facilitate the robustness, we propose a multi-scale convolution-enhanced adaptation method to adapt pre-trained visual foundation models to the VPR task, which introduces the multi-scale local information to further enhance the cross-image correlation-aware representation. Experimental results show that our method outperforms state-of-the-art methods by a large margin with significantly less training time. Our method achieves 94.5% R@1 on Pitts30k using 512-dim global features. The code is released at https://github.com/Lu-Feng/CricaVPR.
1 Introduction
Figure 1: The Recall@1 and descriptors dimensionality comparison of different SOTA methods on Pitts30k. The GCL, NetVLAD, SFRS, and CricaVPR (Ours) all use PCA for dimensionality reduction. Our method achieves significantly higher Recall@1 than other methods with 512-dim compact global features.
Visual place recognition (VPR), also known as visual geo-localization [9, 8], aims at getting the coarse geographical location of an input query image by retrieving the most similar place image from a geo-tagged database. VPR has wide applications in augmented reality [45], mobile robot localization [66], and so on. However, there are three key challenges in VPR: condition (e.g., lighting, weather, and season) variations, viewpoint variations, and perceptual aliasing [40] (difficult to distinguish highly similar images taken from different places). Addressing these challenges at the same time is a hard nut to crack, especially for methods that use only global features.
VPR is typically addressed as an image retrieval problem [12]. The place images are represented using global features and the similarity search is implemented in this feature space to return the matched place image. The global features are usually derived through the aggregation or pooling of local features, employing methods such as NetVLAD [5] or GeM [52] pooling. Such compact features are suitable for large-scale VPR. However, they lack robustness in challenging environments and are often susceptible to perceptual aliasing. An approach to improve robustness is to perform re-ranking by matching local features [26, 61], which incurs huge overhead in both runtime and memory footprint, i.e., not suitable for large-scale VPR. One problem that has been neglected is that existing methods produce the feature of an image only using this image itself (without cross-image interaction), which does not explicitly consider cross-image variations. To address this, our method attempts to use the cross-image variations as a cue to guide the representation learning and harvest useful information from other images when producing the feature of an image, making the output feature condition-invariant, viewpoint-invariant, and capable of addressing perceptual aliasing.
Moreover, the recent visual foundation models [53, 71, 48] have achieved powerful performance. However, due to the particularity of the VPR task, directly using the pre-trained foundation model will encounter some problems. For example, the image features produced using pre-trained models tend to ignore some discriminative backgrounds, and are susceptible to interference from dynamic foregrounds (see Fig. 6 in experiments). Fine-tuning the model on VPR datasets can address this but tends to hurt the previously learned ability, i.e., catastrophic forgetting [19]. A promising way is to exploit parameter-efficient transfer learning (PETL) [28, 29]. However, the discriminative landmarks that need attention in VPR often occupy local regions of uncertain size in images, and most existing PETL methods use language-oriented adaptation modules to adapt the transformer model and lack the image-related (multi-scale) local priors for visual tasks (especially for VPR). This raises the need to develop a new adaptation method to introduce multi-scale local priors to the foundation model for VPR. Taking multi-scale variations into account is also an important complement to our cross-image correlation awareness.
In this paper, we propose a novel method to learn Cross-image correlation-aware representation for VPR, abbreviated as CricaVPR. Our method first uses a backbone with the pooling module to yield initial feature representations. Then we use the cross-image encoder equipped with self-attention to calculate the correlation between multiple image representations within a batch to get final features. The images in a batch can be taken in the same place under different conditions (e.g. lighting) or from different viewpoints, or even captured from different places. This process allows each feature to enhance itself with useful information from others, thus producing condition-robust, viewpoint-robust, and discriminative representations. Meanwhile, we use the visual foundation model as the backbone in our architecture, and design a multi-scale convolution-enhanced adaptation method, in which we freeze the pre-trained foundation models and insert a few trainable lightweight adapters, to introduce the multi-scale local prior knowledge and adapt the foundation model for the VPR task.
Our work brings the following contributions: 1) We propose a cross-image correlation-aware representation method, which uses self-attention to model the correlation between multiple image representations within a batch and make each feature more robust. Our method can also be combined with common aggregation methods to significantly boost performance. 2) We design a parameter-efficient adaptation method to adapt pre-trained models for VPR, in which the proposed multi-scale convolution adapter is used to introduce multi-scale local information to boost performance. 3) Extensive experiments on several VPR benchmark datasets show that our method can outperform the state-of-the-art (SOTA) methods by a large margin (see Fig. 1 for results on Pitts30k) with less training time.
2 Related Work
Visual Place Recognition: The early VPR approaches typically represent place images by global features that are computed using aggregation algorithms, such as Bag of Words [3] and Vector of Locally Aggregated Descriptors (VLAD) [30, 39, 4, 58, 35], to aggregate the hand-crafted descriptors like SURF [6, 16]. Then these methods perform a nearest neighbor search in the global feature space over the database to get the most similar images. With the significant success of deep learning on various computer vision tasks, most recently VPR methods [55, 5, 32, 14, 15, 47, 21, 22, 65, 70, 38, 1, 2, 23, 8, 10] have employed a variety of deep features to represent place images for boosting performance. Likewise, the aggregation algorithm has also been changed into a differentiable module to embed neural networks for end-to-end training [5, 27, 51]. However, most of the global-retrieval-based methods lack robustness in challenging environments and are prone to perceptual aliasing.
Two typical ways to alleviate this issue are to impose temporal consistency constraints and spatial consistency constraints. The former performs image sequence matching (i.e. utilize temporal continuity) [46, 17, 24, 41, 20] to realize robust VPR in challenging environments. The latter is often developed as a two-stage VPR system [25, 26, 7, 61, 72, 42, 43], which searches for top-k candidate images over the database using global features, then performs spatial consistency matching using local features to re-rank candidates. Different from these methods bringing additional constraints, runtime, and memory overhead, our model learns highly robust global representation via cross-image correlation awareness for global-retrieval-based VPR.
Parameter-efficient Transfer Learning: Some recent studies [53, 71, 62, 48] trained the large transformer-based foundation models on huge quantities of data. These models are capable of producing well-generalized feature representation and performing admirably on some common visual tasks. A promising technique for adapting these foundation models to more diverse downstream tasks with only fine-tuning a few (extra) parameters is PETL [28, 37, 29], which is initially proposed in natural language processing to address the catastrophic forgetting issue [19] and reduce training costs. Training the inserted task-specific adapters [28] while keeping the pre-trained foundation models frozen is one of the commonly used PETL methods, and we follow it in our work. There are multiple adapter-based methods [67, 31, 13, 49, 34, 50, 68] have been proposed to address a wide range of visual tasks. A closely related work to ours is Convpass [31], which used convolutional bypasses in ViT as adaptation modules to introduce image-related local inductive biases and avoid performance degradation in downstream fine-tuning. However, our work designs a multi-scale convolution adapter to learn more proper local information to improve the performance on the VPR task.
3 Methodology
Our method involves the Vision Transformer (ViT) model and the self-attention mechanism. So we first briefly review them in this section. Then we propose the cross-image correlation-aware representation method to describe place images. Finally, we present the multi-scale convolution-enhanced adaptation method to adapt the foundation model for VPR and the training strategy for fine-tuning.
3.1 Preliminary
The ViT model [18] and its variants have been applied for many computer vision tasks including VPR mainly due to its superior performance in modeling long-range dependencies. To process an input image with ViT, the image is initially divided into $N$ non-overlapping patches, which are then linearly projected into D-dim patch embeddings $x_{p}\in\mathcal{R}^{N\times D}$. Meanwhile, a learnable [class] token is prepended to $x_{p}$ to form $x_{0}=[x_{class};x_{p}]\in\mathcal{R}^{(N+1)\times D}$. To preserve the original positional information of each patch token, the corresponding positional embeddings are added to $x_{0}$ to get $z_{0}$, which is fed into a series of transformer encoder layers to yield the feature representation. A standard transformer encoder layer consists of three main components: the multi-head attention (MHA) layer, the MLP layer, and the LayerNormalization (LN) layer. The forward process of input $z_{l-1}$ passing through a transformer encoder layer to yield the output $z_{l}$ can be formulated as
$z^{\prime}_{l}$ $=\text{MHA}\left(\text{LN}\left(z_{l-1}\right)\right)+z_{l-1},$ (1)
$z_{l}$ $=\text{MLP}\left(\text{LN}\left(z^{\prime}_{l}\right)\right)+z^{\prime}_{l}.$
The MLP layer is made up of two fully connected layers, which are mainly used for feature nonlinearization and dimension conversion. Here we briefly overview the process of calculating the correlation and attention in the MHA layer. The input $X\in\mathcal{R}^{N\times D}$ is first linearly transformed to produce the queries $Q$, keys $K$, and values $V$. Then the attention among $Q,K$ and $V$ is computed using the Scaled Dot-Product Attention [60], denoted as
$\mathit{Attn}(Q,K,V)=\mathit{Softmax}\left(QK^{\top}/\sqrt{d}\right)V.$ (2)
The MHA first splits the queries, keys, and values into $h$ parts and performs attention for each part in parallel. The outputs of $h$ attention heads are then concatenated. The $N$ elements in the input sequence $X$ are correlated in this process. Specifically, we regard each element as a query, and the remaining elements as keys and values. For each query, we calculate the attention score between it and each key, establishing correlations between every query and all keys. These scores are then multiplied with their corresponding values, modeling dependencies among the $N$ elements.
There are two ways to yield global representations of places using the output of ViT. The first is to directly use the output class token as a global feature. The second is to reshape the output patch tokens as a feature map (similar to the output of CNN) to restore spatial position, and use the aggregation/pooling method (e.g. GeM [52]) to process it as a global feature. Both class token and GeM pooling are used to produce the place representation in our work.
3.2 Cross-image Correlation-aware Place Representation
Figure 2: The example of partial images in a batch. (a), (b), and (c) are taken from the same place with different conditions (seasons) and viewpoints. (d), (e), and (c) are captured from different places, but (d) is similar to (c). When the model produces the features of (c), it can harvest relevant information from other images to yield a better representation.
Figure 3: The pipeline to produce the proposed cross-image correlation-aware representation. The cross-image encoder is the core component for modeling correlations between different image features in a batch. Note that we are correlating the $i$-th regional features of all images in a batch, not all regional features of an image.
The methods based on neural networks have dominated the VPR area over the past decade. These methods commonly produce the deep feature representation of an image with only this image itself. Such features often lack robustness in challenging environments and are incapable of addressing the perceptual aliasing issue. In this work, we present a simple and effective solution to this problem. We attempt to correlate the features of place images in a batch, so that each image representation can harvest useful information from the other image representations to enhance its own robustness. More specifically, there may be images from the same place but taken from different viewpoints or under different conditions, or images from different places that look similar (or not) in a batch, as shown in Fig. 2. On the one hand, image representations from the same place with different perspectives and conditions can improve the viewpoint invariance and condition invariance of each other after the correlated encoding. On the other hand, image representations from different places also promote each other to produce discriminative features. As a result, our method can provide highly robust global representations to deal with viewpoint changes, condition changes, and perceptual aliasing.
We design the pipeline to produce desired global features as shown in Fig. 3. The output patch tokens of a batch of images from the ViT backbone are reshaped as the $B\times W\times H\times D$-dim (i.e., batchbatch sizesize × weightweight × heightheight × tokentoken dimensiondimension) feature maps. We first use the spatial pyramid [36] to produce initial feature representations. The feature maps are split at three levels ($\times1$, $\times2$, and $\times3$). Then we use GeM pooling to process local (patch) features within the divided regions and get a total of 14 regional features of each image. Since the first level is a global aggregation, we directly use the class token to replace the GeM feature in this level for better performance. Next is the most critical step. We treat the $i$-th regional features of all images in a batch as a sequence of embedding vectors $f_{i}$, that is
$f_{i}=\{f_{i}^{1},f_{i}^{2},...,f_{i}^{B}\}\quad i\in\{1,2,...,14\},$ (3)
and feed the 14 sequences of embedding vectors into the cross-image encoder to model the correlation between the $i$-th regional features of all images in a batch. Instead of directly using the self-attention layer, the cross-image encoder is structured using two transformer encoder layers that also include the subsequent MLP layer and the skip connection for stable training and better performance. That is, we apply Eq. 1 to process each $f_{i}$ to correlate image representations in a batch. The 14 output regional features of each image are sequentially concatenated (i.e., flattened) as the final global representation of the image.
It should be noted that the retrieval process of our method is the same as the common global-retrieval-based method. That is, it uses the global feature of a single image for retrieval, which does not bring additional restrictions. Besides, we choose the spatial pyramid to produce the initial feature in order to subsequently correlate images at different scales, and the final sequential concatenation of regional features also introduces spatial position information to the global representations. In fact, our method can also use other aggregation methods to yield initial features, and boost the performance of these methods.
3.3 Multi-scale Convolution-enhanced Adaptation
Our work adapts the distilled DINOv2 [48] as the backbone (i.e. the adapted DINOv2/ViT backbone in Fig. 3), which is based on ViT-B/14. The DINOv2 work trains the ViT model on the large-scale curated LVD-142M dataset with the self-supervised strategy, and can provide powerful visual features to achieve promising performance on some common tasks without any fine-tuning. AnyLoc [33] is a VPR work that uses pre-trained DINOv2 without fine-tuning. However, there exists a gap between the tasks of model pre-training and VPR due to the inherent difference in training objectives and data. Directly using such a pre-trained model in VPR cannot fully unleash its powerful capability.
The adapter-based parameter-efficient transfer learning [28] provides an effective way to adapt foundation models for downstream tasks, which freezes the pre-trained model and only fine-tunes the added lightweight adapter. The vanilla adapter is a bottleneck module consisting of a down-projection (fully connection) layer, an up-projection layer, and a non-linearity (activation) layer in the middle. The Convpass work [31] applies convolution layers to introduce image-related local inductive biases into models. However, we found that improper local priors provided by Convpass risk reducing performance in VPR. Inspired by the inception module in GoogLeNet [56], we design our multi-scale convolution (MulConv) adapter as shown in Fig. 4 (b).
Different from the vanilla adapter, our MulConv adapter adds a MulConv module between the (ReLU) activation layer and the up-projection layer. Meanwhile, there is a skip connection in parallel to the MulConv module. The MulConv module consists of three parallel convolutional paths of different scales ($\times1$, $\times3$, $\times5$). The $\times1$ convolution is also used before the $\times3$ and $\times5$ convolutions to reduce channel dimension. This design and the bottleneck structure of the adapter make our MulConv adapter still lightweight. The outputs of the three convolutional paths are concatenated to form the output of the MulConv module. Finally, the MulConv adapter is added in parallel to the MLP layer in each transformer block (i.e. transformer encoder layer) of the ViT backbone to achieve multi-scale convolution-enhanced adaptation, which can introduce proper (multi-scale) local priors to the model and improve performance for VPR. So the computation of each adapted transformer block can be denoted as
$z^{\prime}_{l}$ $=\text{MHA}\left(\text{LN}\left(z_{l-1}\right)\right)+z_{l-1},$ (4)
$z_{l}$ $=\text{MLP}\left(\text{LN}\left(z^{\prime}_{l}\right)\right)+s\cdot\text{Adapter}\left(\text{LN}\left(z^{\prime}_{l}\right)\right)+z^{\prime}_{l}.$
Figure 4: Illustration of our multi-scale convolution-enhanced adaptation. (a) is a standard transformer block. (b) is the MulConv adapter. We add the MulConv adapter in parallel to the MLP layer in each transformer block to achieve our adaptation as in (c).
3.4 Training Strategy
Different from the common VPR works training models with weak supervision, we train our model on the GSV-Cities [1] dataset with full supervision. This dataset contains 560k images captured at 67k places with highly accurate labels. We follow the standard framework of this dataset and use the multi-similarity (MS) loss [63] with online hard mining strategy for training, as in the GSV-Cities [1] and MixVPR [2] works. The MS loss is computed as
$\begin{split}\mathcal{L}_{MS}=\frac{1}{m}\sum_{q=1}^{m}\bigg\{\frac{1}{\alpha}{\log\big[1+\sum_{p\in\mathcal{P}_{q}}e^{-\alpha(S_{qp}-\lambda)}}\big]\\ +\frac{1}{\beta}{\log\big[1+\sum_{n\in\mathcal{N}_{q}}e^{\beta(S_{qn}-\lambda)}\big]}\bigg\},\end{split}$ (5)
where for each query (anchor) image $I_{q}$ in a batch, $\mathcal{P}_{q}$ is the set of indices $\{p\}$ that correspond to the positive samples for $I_{q}$, and $\mathcal{N}_{q}$ is the set of indices $\{n\}$ that correspond to the negative samples for $I_{q}$. $S_{qp}$ and $S_{qn}$ are the cosine similarities of a positive pair $\{I_{q},I_{p}\}$ and a negative pair $\{I_{q},I_{n}\}$. $\alpha$, $\beta$ and $\lambda$ are three set constants (hyperparameters).
4 Experiments
4.1 Datasets and Performance Evaluation
Dataset Description Number
Database Queries
Pitts30k urban, panorama 10,000 6,816
MSLS-val urban, suburban 18,871 740
MSLS-challenge long-term 38,770 27,092
Tokyo24/7 urban, day/night 75,984 315
Nordland natural, seasonal 27,592 27,592
SVOX cross-domain 17,166 4,356
AmsterTime very long-term 1,231 1,231
Table 1: Summary of the test datasets in experiments.
The experiments are conducted on several VPR benchmark datasets. These datasets exhibit viewpoint changes, condition changes, and the perceptual aliasing issue. Table 1 summarizes the key information of them. Pitts30k [57] contains a total of 30k database images in train, validation, and test sets. We use the test set to evaluate models. MSLS [64] consists of more than 1.6 million images captured in urban, suburban, and natural scenes from 30 cities on six continents over 7 years. We evaluate models on both MSLS-val and MSLS-challenge sets. The latter is an online test set without publicly released labels. Tokyo24/7 [58] contains about 76k reference images and 315 query images collected in urban scenes, and shows severe illumination (day/night) changes. We also use three challenging datasets: Nordland (with seasonal changes) [10], SVOX (cross-domain dataset) [11], and AmsterTime (with very long-term changes) [69]. More details are in Supplementary (Suppl.) Material.
The Recall@N (R@N) metric is used in our experiments to evaluate recognition performance. It is the percentage of queries for which at least one of the N retrieved database images is taken within a threshold of ground truth. We set the threshold to 25 meters and 40 for MSLS, 25 meters for Pitts30k, Tokyo24/7, and SVOX, $\pm 10$ frames for Nordland, unique counterpart for AmsterTime, following common evaluation procedures [64, 57, 58].
Method Dim Pitts30k Tokyo24/7 MSLS-val MSLS-challenge
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
NetVLAD [5] 32768 81.9 91.2 93.7 60.6 68.9 74.6 53.1 66.5 71.1 35.1 47.4 51.7
SFRS [23] 4096 89.4 94.7 95.9 81.0 88.3 92.4 69.2 80.3 83.1 41.6 52.0 56.3
Patch-NetVLAD [26] / 88.7 94.5 95.9 86.0 88.6 90.5 79.5 86.2 87.7 48.1 57.6 60.5
TransVPR [61] / 89.0 94.9 96.2 79.0 82.2 85.1 86.8 91.2 92.4 63.9 74.0 77.5
CosPlace [8] 512 88.4 94.5 95.7 81.9 90.2 92.7 82.8 89.7 92.0 61.4 72.0 76.6
GCL [38] 2048 80.7 91.5 93.9 69.5 81.0 85.1 79.5 88.1 90.1 57.9 70.7 75.7
MixVPR [2] 4096 91.5 95.5 96.3 85.1 91.7 94.3 88.0 92.7 94.6 64.0 75.9 80.6
EigenPlaces [10] 2048 92.5 96.8 97.6 93.0 96.2 97.5 89.1 93.8 95.0 67.4 77.1 81.7
CricaVPR (ours) 4096 94.9 97.3 98.2 93.0 97.5 98.1 90.0 95.4 96.4 69.0 82.1 85.7
Table 2: Comparison to state-of-the-art methods on benchmark datasets. The best is highlighted in bold and the second is underlined.
Figure 5: Qualitative results. These four challenging examples show severe viewpoint changes and condition changes. The proposed CricaVPR successfully yields the right results, while other methods return incorrect images. In each example, there are methods to return similar images from different places (i.e., incorrect) due to perceptual aliasing. In the second example, the query image is taken at night, causing all the other methods to return night images but from different places (i.e. wrong). However, our method returns an image taken during the day at the same place (i.e. correct).
Method Nordland Amster SVOX SVOX
Time -Night -Rain
SFRS [23] 16.0 29.7 28.6 69.7
CosPlace [8] 58.5 38.7 44.8 85.2
MixVPR [2] 76.2 40.2 64.4 91.5
EigenPlaces [10] 71.2 48.9 58.9 90.0
CricaVPR (ours) 90.7 64.7 85.1 95.0
Table 3: Comparison (R@1) to SOTA methods on more challenging datasets. More results are in Suppl. Material.
4.2 Implementation Details
We fine-tune our model on two NVIDIA GeForce RTX 3090 GPUs using PyTorch. The resolution of the input image is 224×224 and the token dimension of the backbone (ViT-B/14) is 768. Our model outputs the 14×768-dim original global features, and we use PCA for dimensionality reduction. The bottleneck ratio of our MulConv adapters is set to 0.5, so the input dimension of the three convolutional paths is 384. The 1×1 convolution before the 3×3 and 5×5 convolution reduces the channel to 24. The output dimensions of the three convolutional paths are 192, 96, and 96. We set the hyperparameters $\alpha=1,\beta=50,\lambda=0$ in Eq. 5 and margin = 0.1 in online mining, as in GSV-Cities [1]. We fine-tune the model using the Adam optimizer with the initial learning rate set as 0.0001 and multiplied by 0.5 after every 3 epochs. A training batch contains 72 places with 4 images each (i.e. 288 images). Training is implemented until the R@5 on Pitts30k does not improve for 3 epochs. An inference batch contains 8 images for Pitts30k and 16 images for others.
4.3 Comparisons with State-of-the-Art Methods
In this section, we compare our CricaVPR with several SOTA VPR methods, mainly including six global-retrieval-based methods: NetVLAD [5], SFRS [23], CosPlace [8], GCL [38], MixVPR [2] and EigenPlaces [10]. Note that our work uses the same training dataset as MixVPR, i.e., GSV-Cities. Meanwhile, CosPlace and EigenPlaces are trained on individually constructed extra large-scale datasets, i.e., SF-XL. Both MixVPR and EigenPlaces are the latest works and represent the SOTA performance of the VPR methods based on global feature retrieval. Additionally, we also compare our approach with two excellent two-stage VPR methods (Patch-NetVLAD [26] and TransVPR [61]), which require time-consuming re-ranking using local features. The details of these methods are in Suppl. Material. Table 2 shows the quantitative results on Pitts30k, Tokyo24/7, and MSLS. Our CricaVPR method achieves the best R@1/R@5/R@10 on all datasets.
MixVPR, EigenPlaces, and our CricaVPR all achieve excellent performance on these datasets. Especially on Pitts30k, which shows significant viewpoint changes but no drastic condition changes, EigenPlaces achieves 92.5% R@1. This indicates that the challenge posed by viewpoint changes has been effectively addressed by existing methods (i.e., EigenPlaces and MixVPR). However, our method continues to improve performance on Pitts30k, achieving an impressive 94.9% R@1. This improvement primarily stems from the powerful ability of our method to produce more discriminative global representations to differentiate similar images from different places, i.e., address perceptual aliasing. The MSLS dataset is more challenging as it shows severe condition variations and includes some suburban or natural scene images lacking landmarks and prone to perceptual aliasing. Nevertheless, our method achieves 95.4% R@5 on MSLS-val and 82.1% R@5 on MSLS-challenge, showing significant advantages over other global-retrieval-based methods and two-stage methods.
Fig. 5 qualitatively demonstrates the superior performance of our method in some extreme environments. These challenging examples include drastic condition changes, viewpoint changes, or only small regions in the images showing discriminative objects. In these examples, other methods either get similar images but from different places (i.e. suffer from perceptual aliasing), or retrieve places that are close in geographical distance but still out of the set threshold, that is, they fail to retrieve the correct results. Our approach shows high robustness against these challenges.
To further evaluate the performance of our method in extreme scenarios, we conduct experiments on three challenging datasets: Nordland, which exhibits seasonal changes; AmsterTime, which spans a very long time period; and SVOX, which shows extreme illumination and weather variations. The results, as shown in Table 3, demonstrate the significant superiority of our method compared to other SOTA methods. Our CricaVPR outperforms all other SOTA methods with 14.5%, 15.8%, and 20.7% absolute R@1 improvements on Nordland, AmsterTime, and SVOX-Night, respectively. This further highlights that our global representation is highly robust.
Moreover, our model only costs 3.5 hours for training, which is significantly less than the full-day time used by CosPlace/EigenPlaces. The training epochs of our model (10 epochs) are also less than MixVPR (30 epochs) using the same dataset. More details are in Suppl. Material.
4.4 Ablation Study
We perform a series of ablation experiments to validate the effectiveness of the proposed components in our method. All ablated methods no longer use PCA for dimensionality reduction by default. We will conduct separate experiments to show the impact of feature dimensions on the results.
Ablated versions Pitts30k Tokyo24/7 MSLS-val
R@1 R@5 R@1 R@5 R@1 R@5
FrozenDINOv2-SPM 74.8 90.1 49.8 67.0 45.4 60.7
AdaptGeM 87.1 94.0 70.2 85.4 78.4 87.8
AdaptSPMG 87.8 94.1 72.1 85.1 78.0 88.4
AdaptSPM 90.6 95.9 85.1 93.3 85.5 93.2
AdaptGeM+Crica 93.9 97.2 87.6 93.3 86.1 93.4
AdaptSPMG+Crica 94.3 97.3 93.7 96.5 89.7 95.3
AdaptSPM+Crica 94.8 97.4 93.0 97.1 89.9 95.4
Table 4: Ablation on cross-image awareness. The “+Crica” represents the addition of our cross-image correlation awareness to get the final global feature. The “SPM” represents our spatial pyramid model representation, while “SPMG” is the spatial pyramid model solely based on GeM. Except for the DINOv2-GeM and DINOv2-SPM that directly use an untuned backbone (as baseline), all other versions use our adaptation method for fine-tuning.
Ablation on cross-image correlation awareness. The cross-image correlation awareness achieved by the cross-image encoder after backbone is the most important module in our method. We compare the performance of the three kinds of global features with or without across-image awareness. These features are GeM, the spatial pyramid model representation solely based on GeM (SPMG), and our spatial pyramid model representation (SPM) which uses both class token and GeM. The results are shown in Table 4. After incorporating the proposed cross-image correlation awareness (Crica), all three features achieve significant performance improvements. Due to the already impressive performance of the SPM feature after our model adaptation (AdaptSPM), the improvement provided by Crica on this feature is not as pronounced as on the GeM and SPMG features. Nevertheless, AdaptSPM+Crica still resulted in 4.2%, 7.9%, and 4.4% absolute R@1 improvements over the AdaptSPM feature on Pitts30k, Tokyo24/7, and MSLS-val, respectively. Moreover, AdaptGeM+Crica achieves an impressive 17.4% absolute R@1 improvement over AdaptGeM on Tokyo24/7. With the combined effect of our Crica and model adaptation, our method achieves nearly 2× higher R@1 on Tokyo24/7 and MSLS-val compared to the direct use of frozen DINOv2 with the SPM representation (FrozenDINOv2-SPM). In the next subsection, we will further validate our model adaptation.
Ablation on adaptation. We first use only the GeM features alone (without cross-image awareness) to demonstrate the performance improvement achieved by our adaptation method. As shown in Table 5, MulConvAdapter-GeM using our adaptation achieves a significant improvement over FrozenDINOv2-GeM. Especially on MSLS-val, which has more dynamic interference, our adaptation achieves nearly 2× higher R@1. Fig. 6 vividly illustrates the underlying reasons. The adapted DINOv2, in contrast to the pre-trained DINOv2, exhibits a greater ability to attend to objects that are relevant for place recognition, with more attention given to more important objects. Table 5 also shows the performance of different fine-tuning methods when using the proposed global features (i.e. the SPM feature with cross-image correlation awareness). The FullTunedDINOv2 achieves a notable improvement over FrozenDINOv2 on Pitts30k and MSLS-val. However, because our training data has no night images like those in Tokyo24/7, FullTunedDINOv2 performs worse than FrozenDINOv2 on Tokyo24/7, i.e., it suffers from catastrophic forgetting. This indicates the necessity of parameter-efficient fine-tuning (using an adapter). Besides, ConvAdapter (as in Convpass [31]) uses 3×3 convolution to introduce local inductive biases into the model. However, it brings inappropriate local priors for VPR and results in performance degradation compared to VanillaAdapter. Our method (MulConvAdapter) uses multi-scale convolution to introduce more proper local information and thus achieves the best performance.
Figure 6: The output feature map (attention) visualizations of pre-trained DINOv2 and adapted DINOv2. The regions attended to by pre-trained DINOv2 have no relevance to place recognition. However, adapted DINOv2 focuses on discriminative areas for VPR. Buildings that remain relatively unchanged over time receive the highest attention. Vegetation that is not expected to change in the short term receives moderate attention. Non-discriminative elements such as the sky and ground, as well as dynamic vehicles, are ignored.
Ablated versions Pitts30k Tokyo24/7 MSLS-val
R@1 R@5 R@1 R@5 R@1 R@5
FrozenDINOv2-GeM 79.2 90.1 65.4 83.8 40.8 51.5
MulConvAdapter-GeM 87.1 94.0 70.2 85.4 78.4 87.8
FrozenDINOv2 79.2 90.1 80.0 89.8 58.8 71.2
FullTunedDINOv2 94.1 96.6 76.8 88.3 86.2 93.2
VanillaAdapter 94.6 97.4 92.7 96.5 89.2 95.5
ConvAdapter 93.8 96.9 92.7 95.9 88.0 94.2
MulConvAdapter 94.8 97.4 93.0 97.1 89.9 95.4
Table 5: Ablation on adaptation. Except for the versions with the ”-GeM” suffix, which utilize GeM features, all other versions use our spatial pyramid representation with the proposed cross-image awareness to yield global features. FrozenDINOv2 and FullTunedDINOv2 represent the use of frozen and fully fine-tuned DINOv2 as backbones, respectively. VanillaAdapter, ConvAdapter, and MulConvAdapter represent the use of a vanilla adapter, 3x3 convolution adapter, and our proposed multi-scale convolution adapter to adapt DINOv2 as the backbone, respectively.
Ablation on dimensions of our descriptor. Intuitively, higher-dimensional descriptors have the ability to provide more information, thus offering the potential to achieve better performance. However, removing redundant or noisy information often leads to performance improvements as well. We analyze the impact of descriptor dimensionality on performance, and the results are shown in Table 6. Our method gets the best performance when using PCA to reduce the descriptor dimension to 4096-dim, so it is the default dimensionality we recommend. Furthermore, we continue to reduce the dimensionality to observe the point at which performance starts to noticeably decline on each dataset. For Pitts30k, the 512-dim descriptor still achieves an impressive 94.5% R@1, with no significant decrease compared to the 4096-dim descriptor. However, using the 512-dim descriptor on the other two datasets results in an obvious performance drop. This is mainly due to the drastic condition changes and perceptual aliasing issue in these datasets, requiring higher-dimensional descriptors to provide sufficient information to distinguish places. When there is a pressing need for low-dimensional descriptors, we suggest using the 512-dim descriptor for Pitts30k, while selecting the 1024-dim or 2048-dim descriptor for Tokyo24/7 and MSLS.
Dim Pitts30k Tokyo24/7 MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
512 94.5 97.1 98.0 84.4 93.7 95.9 85.3 93.4 94.5
1024 94.8 97.3 98.1 91.4 96.8 97.8 87.7 94.3 95.3
2048 94.8 97.4 98.2 92.4 96.8 97.8 89.2 95.1 96.1
4096 94.9 97.3 98.2 93.0 97.5 98.1 90.0 95.4 96.4
10752 94.8 97.4 98.1 93.0 97.1 97.8 89.9 95.4 96.2
Table 6: Ablation on dimensions of our descriptor. The original output dimension is 10752.
5 Conclusions
In this paper, we presented CricaVPR, a robust global representation method with cross-image correlation awareness for VPR. Our method leverages the cross-image encoder equipped with self-attention to establish the correlation among multiple images within a batch, enabling the model to harvest useful information from other images while generating the feature representation of an image. This makes the produced global features condition-invariant, viewpoint-invariant, and capable of addressing perceptual aliasing. Furthermore, we proposed a multi-scale convolution-enhanced adaptation method to introduce proper local information and effectively unleash the capability of the pre-trained foundation model for VPR. Experimental results on several VPR benchmark datasets demonstrate that our CricaVPR can provide a robust global representation to address various challenges in VPR and outperforms SOTA methods by a significant margin.
Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant (KJZD20230923115106012), Shenzhen Key Laboratory (ZDSYS20210623092001004), Beijing Key Lab of Networked Multimedia, and the Project of Peng Cheng Laboratory (PCL2023A08).
Supplementary Material
6 Overview
This supplementary material provides the following additional content about experimental results and analysis:
7. Training Epoch, Training Time & Tunable Parameters
8. Visualizations of Place Features Using t-SNE
9. Additional Results on Challenging Datasets
10. Additional Ablations on Cross-image Encoder
11. Effects of Training Batch Size
12. Additional Ablations on MulConvAdapter
13. Effects of Adaptation on the Used SPM Feature
14. Datasets Details
15. Compared Methods Details
16. Additional Qualitative Results and Failure Cases
7 Training Epoch, Training Time, and Tunable Parameters
In the main paper, we have described that our model only trained for 10 epochs on the GSV-Cities [1] dataset using 3.5 hours. This is less than the full-day training time used by CosPlace/EigenPlaces, and less than the 30 training epochs used by MixVPR (also trained on GSV-Cities). To further investigate the data efficiency of our method, we reduce the training epoch (and training data), and the yielded results are shown in Table 7. When the model is trained with only 10% of the training data for 1 epoch (i.e., 0.1 epoch), our method achieves better performance than previous methods (except EigenPlaces), e.g., CosPlace and MixVPR, on Pitts30k and Tokyo24/7. The training time used is only 0.038h (i.e., 2.3min). The advantages of our method in data efficiency are mainly due to the fact that the adapter-based method maintains the powerful representation ability of the pre-trained foundation model, while our proposed cross-image encoder is an easy-to-train module. Our model has little risk of over-fitting even when trained with small amounts of data.
We provide detailed model parameters as shown in Table 8 (using CosPlace as baseline). Since we use an adapter-based parameter-efficient fine-tuning method to train our model, the tunable part of our model only contains the adapter inserted into the backbone and the cross-image encoder after the backbone. The number of tunable parameters of our model is 20.2M, which is only about 1/5 of the full fine-tuning DINOv2 (with the cross-image encoder). This is also less than that of CosPlace using ResNet50 to produce 2048-dim features (including 26.3M tunable parameters in the model and 582.8M tunable parameters of classifiers).
Epoch Training Pitts30k Tokyo24/7 MSLS-val
time (h) R@1 R@5 R@1 R@5 R@1 R@5
10 3.5 94.8 97.4 93.0 97.1 89.9 95.4
5 1.8 94.0 97.2 93.3 96.2 89.1 95.3
1 0.36 93.3 96.7 92.7 95.9 85.4 93.8
0.1 0.038 92.5 96.5 88.9 96.2 79.1 88.4
Table 7: The results of our CricaVPR with different training epochs.
Method Total Backbone Adapter Others Tunable
CosPlace-V - 14.7 0 0.3 7.3+145.7
CosPlace-R - 23.5 0 4.2 26.3+582.8
FullTuning 97.6 86.6 0 11.0 97.6
Ours 106.8 86.6+9.2 9.2 11.0 20.2
Table 8: The number of parameters (M) in models. We mainly focus on tunable parameters (the last column). “CosPlace-V” and “CosPlace-R” represent the CosPlace methods using VGG16 and ResNet50 to produce 512-dim and 2048-dim features, respectively. Taking CosPlace-V as an example, since it adds multiple classifiers (for multiple groups of training data) after the model during training, the tunable parameters contain the parameters of the trainable part in the model (7.3M) and all classifiers (145.7M). “FullTuning” represents full fine-tuning of the DINOv2 backbone (including our cross-image encoder) without the adapter. The “Others” in the table are the aggregation module for CosPlace, the cross-image encoder for FullTuning and Ours (the parameters of GeM pooling are so few that they can be ignored).
Figure 7: Visualizations of place features in 2-dimensional space using t-SNE. We use the features of 432 images from 36 different places (i.e. 36 categories) for visualization. (a), (b), and (c) are the results of pre-trained DINOv2, adapted DINOv2 (using our MulConv adapter), and our complete model (using MulConv adapter and cross-image encoder), respectively. Note that the positions of two points in (b) are improper, that is, the corresponding feature representation will suffer perceptual aliasing.
Method Nordland AmsterTime SVOX-Night SVOX-Rain SVOX-Sun
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
SFRS [23] 16.0 24.1 28.7 29.7 48.5 55.6 28.6 40.6 46.4 69.7 81.5 84.6 54.8 68.3 74.1
CosPlace [8] 58.5 73.7 79.4 38.7 61.3 67.3 44.8 63.5 70.0 85.2 91.7 93.8 67.3 79.2 83.8
MixVPR [2] 76.2 86.9 90.3 40.2 59.1 64.6 64.4 79.2 83.1 91.5 97.2 98.1 84.8 93.2 94.7
EigenPlaces [10] 71.2 83.8 88.1 48.9 69.5 76.0 58.9 76.9 82.6 90.0 96.4 98.0 86.4 95.0 96.4
CricaVPR (ours) 90.7 96.3 97.6 64.7 82.8 87.5 85.1 95.0 96.7 95.0 98.2 98.7 93.7 98.4 98.6
Table 9: Comparison to SOTA methods on challenging datasets. The best is highlighted in bold and the second is underlined. As in the main paper when comparing with SOTA methods, we employ PCA to reduce the descriptor dimension of our method to 4096-dim.
8 Visualizations of Place Features using t-SNE
In this section, we use the t-SNE [59] method to map our place features to 2-dimensional space and visualize their distribution. We employ pre-trained DINOv2, adapted DINOv2 (with our MulConv adapter), and our entire network (with our MulConv adapter and cross-image encoder) to extract features of 432 images from 36 different places (12 images per place). There exist variations in viewpoints and conditions among the 12 images of the same place. Fig. 7 illustrates the visualization results. It can be observed that some features of different places, which are extracted by pre-trained DINOv2, are not well separated. This demonstrates the limited discriminability of place features extracted by pre-trained DINOv2. However, after performing our adaptation, the adapted DINOv2 successfully distinguishes most of the places, while a few places are still not well distinguished. By applying both our adaptation and the cross-image encoder, our proposed model effectively clusters image features of the same place and separates features of different places, i.e., pulling the features of the same place closer together and pushing the features of different places farther apart. This clearly demonstrates the efficacy of our approach in addressing the challenge of perceptual aliasing.
It is worth mentioning that this visualization method commonly used in classification tasks has rarely been used in previous VPR works. We can use it thanks to the recently proposed GSV-Cities dataset [1] (and the SF-XL dataset [8]) that split place images into a finite number of categories.
9 Additional Results on Challenging Datasets
The main paper has presented the R@1 results of our method compared to state-of-the-art (SOTA) methods on three challenging datasets, i.e., Nordland, AmsterTime, and SVOX (SVOX-Night, SVOX-Rain). Here, we provide the complete R@1/R@5/R@10 results as shown in Table 9, complementing another challenging query subset (SVOX-Sun) of the SVOX dataset. Before our method was proposed, MixVPR and EigenPlaces had their own advantages on these challenging datasets, and no method completely outperformed the other methods. However, our proposed CricaVPR achieves better performance compared to all previous methods on these datasets, particularly outperforming other methods by a large margin on Nordland, AmsterTime, and SVOX-Night, which are quite difficult.
10 Additional Ablations on Cross-image Encoder
Cross-image encoder Pitts30k Tokyo24/7 MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
No encoder 90.6 95.9 97.2 85.1 93.3 95.6 85.5 93.2 94.3
Transformer encoder layer $\times$1 92.9 96.6 97.5 92.7 95.2 96.5 89.6 95.7 96.4
Transformer encoder layer $\times$2 94.8 97.4 98.1 93.0 97.1 97.8 89.9 95.4 96.2
Transformer encoder layer $\times$3 94.5 97.4 98.1 93.0 96.2 97.8 88.8 94.7 96.1
Table 10: The results of constructing the cross-image encoder using different numbers of transformer encoder layers.
Batch Size with cross- Pitts30k Tokyo24/7 MSLS-val
(Number of Places) image encoder R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
NPNP = 16 $\times$ 89.5 95.3 96.8 75.6 89.2 91.4 80.9 90.4 92.7
NPNP = 32 $\times$ 89.9 95.2 96.7 81.3 91.1 93.0 83.0 93.1 93.8
NPNP = 64 $\times$ 90.7 95.9 97.5 84.4 94.3 96.5 84.1 92.3 94.2
NPNP = 72 $\times$ 90.6 95.9 97.2 85.1 93.3 95.6 85.5 93.2 94.3
NPNP = 16 94.6 97.0 97.7 87.9 94.9 96.2 84.1 92.6 94.2
NPNP = 32 94.8 97.4 98.0 91.1 94.9 96.8 85.0 93.1 95.1
NPNP = 64 94.9 97.5 98.1 92.4 95.6 97.1 88.1 95.0 95.1
NPNP = 72 94.8 97.4 98.1 93.0 97.1 97.8 89.9 95.4 96.2
Table 11: Results of different training batch sizes, i.e., different numbers of places (4 images per place). NPNP is the abbreviation of “Number of Places”. We provide the results with or without the cross-image encoder.
In the main paper, we have combined the proposed cross-image correlation awareness implemented by our cross-image encoder with three different global representations to demonstrate its effectiveness. In this section, we further compare the performance of constructing the cross-image encoder using different numbers of transformer encoder layers, and the results are shown in Table 10. Compared to not using the cross-image encoder (No encoder), incorporating the cross-image encoder constructed with any number of transformer encoder layers leads to significant performance improvements. However, when only one transformer encoder layer is used, there is still a noticeable performance gap compared to using multiple transformer encoder layers (on Pitts30k and Tokyo24/7), indicating that a single transformer encoder layer alone cannot sufficiently correlate images within a batch. The best performance is achieved when using two transformer encoder layers, which is the recommended configuration.
11 Effects of Training Batch Size
Since our method correlates all images within a batch and utilizes the cross-image variations (including images from the same place and images from different places) as a cue to guide the representation learning in VPR, the training batch size is also a factor that may have an impact on performance. The training dataset GSV-Cities [1] provides 4 images per place by default, and we use different batch sizes, i.e., one batch contains different numbers of places, to train our models. It should be noted that we use the multi-similarity (MS) loss to train the models (same as MixVPR), which inherently leads to a result that a larger batch size is more conducive to providing hard sample pairs to train a robust model. Therefore, we also provide the results obtained at different batch sizes without using the cross-image encoder as a reference. The results are shown in Table 11. Regardless of whether the cross-image encoder is used, the performance degradation caused by the smaller batch size is not obvious on Pitts30k, but is significant on more difficult Tokyo24/7 and MSLS. When using the cross-image encoder, the absolute R@1 drops caused by using the smallest batch size ($NP=16$) compared to the largest batch size ($NP=72$) on Pitts30k, Tokyo24/7, and MSLS-val are 0.2%, 5.1%, and 5.8% respectively. When the cross-image encoder is not used, the absolute R@1 drops caused by that are 1.1%, 9.5%, and 4.6% respectively. This indicates that: 1) More challenging (test) datasets require larger batch size to train a more robust model. 2) Our proposed cross-image encoder, to some extent, reduces the demand for a larger batch size when using the MS loss for training.
12 Additional Ablations on MulConvAdapter
We have verified the effectiveness of the proposed multi-scale convolution adapter (MulConvAdapter) by comparing it with the vanilla adapter and ConvAdapter (i.e., Convpass [31]). To further demonstrate the advantages of MulConvAdapter over adapters using only a single-size convolution kernel, we compare MulConvAdapter with three adapter variants employing three different convolution kernel sizes (1×1, 3×3, and 5×5). To be fair, the three adapters based on a single convolution kernel use skip connection like our MulConvAdapter (the ConvAdapter in the main paper does not), that is, our MulConvAdapter differs from these three adapters only in the convolution kernel. The results are presented in Table 12. Except for our MulConvAdapter, the adapters based on 1×1, 3×3, and 5×5 convolution kernels have advantages in different datasets (and metrics), indicating that it is difficult for an adapter with a single-size convolution kernel to perform well for all place images on the VPR task. In contrast, our MulConvAdapter integrates these three convolution kernels to consistently provide proper local information, thus achieving the best performance.
Conv Pitts30k Tokyo24/7 MSLS-val
Size R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
1×1 94.5 97.2 97.8 91.7 95.9 97.1 88.2 95.3 95.5
3×3 94.3 97.1 97.9 91.7 95.2 96.8 87.6 94.3 95.8
5×5 94.7 97.3 97.8 90.2 94.6 96.8 87.6 95.1 96.4
MulConv 94.8 97.4 98.1 93.0 97.1 97.8 89.9 95.4 96.2
Table 12: The results of convolution-based adapters. “MulConv” is our MulConvAdapter. Note that the single convolution kernel adapter here has one more skip connection than the 3×3 convolution adapter (ConvAdapter) in the main paper.
13 Effects of Adaptation on the Used SPM Feature
In our method, we mainly use the spatial pyramid model (SPM) representation that combines class tokens and GeM. An interesting phenomenon is that when using the frozen DINOv2 as the backbone, the SPM feature (FrozenDINOv2-SPM) performs worse than GeM (FrozenDINOv2-GeM) on Pitts30k and Tokyo24/7 (see Table 13). However, after using our adaptation, Adapt-SPM performs much better than Adapt-GeM. This shows that our adaptation makes the combined class token and GeM features in the SPM representation more compatible.
Ablated versions Pitts30k Tokyo24/7 MSLS-val
R@1 R@5 R@1 R@5 R@1 R@5
FrozenDINOv2-GeM 79.2 90.1 65.4 83.8 40.8 51.5
FrozenDINOv2-SPM 74.8 90.1 49.8 67.0 45.4 60.7
Adapt-GeM 87.1 94.0 70.2 85.4 78.4 87.8
Adapt-SPM 90.6 95.9 85.1 93.3 85.5 93.2
Table 13: The results of the GeM and SPM representation using a frozen DINOv2 or adapted DINOv2 backbone. All results here have been provided in Table 4 and Table 5 of our main paper.
14 Datasets Details
Pitts30k [57] is derived from Google Street View panoramas with GPS labels. It consists of images from 24 different viewpoints for each place in urban scenes, exhibiting significant viewpoint variations, moderate condition variations, and a small number of dynamic objects. For our experiments, we use the Pitts30k test set, which comprises 10k database images and 6,816 query images.
Tokyo24/7 [58] comprises a total of 75,984 database images and 315 query images from urban environments. The query images are selected from a pool of 1,125 images captured from 125 places, each involving 3 different viewpoints and 3 different times of the day. This dataset shows viewpoint variations and significant condition changes, particularly day-night changes.
MSLS (Mapillary Street-Level Sequences) [64] is a large-scale VPR dataset that encompasses more than 1.6 million images captured in urban, suburban, and natural environments across 30 cities spanning six continents. This dataset provides GPS coordinates and compass angles for each image, and shows various changes caused by illumination, weather, season, viewpoint, dynamic objects, and so on. It is divided into three sets: training, public validation (MSLS-val), and withheld test (MSLS-challenge). To ensure comprehensive evaluation, we assess the model on both the MSLS-val and MSLS-challenge sets, as done in previous works [26, 61].
Nordland [54] captures images from a fixed viewpoint in the front of a train in four seasons. This dataset exhibits significant variations in conditions such as season and lighting, without viewpoint changes. Its images primarily depict suburban and natural environments, and the ground truth information is provided through frame-level correspondence. Following previous works [10], we extract images at 1FPS, and use the winter images as queries and the summer images for reference (i.e. database).
AmsterTime [69] contains more than a thousand query-reference image pairs captured from Amsterdam. Each pair consists of a grayscale historical image as the query and a contemporary image from the same place (identified by human experts) as the reference. The dataset involves diverse domain variations in viewpoints, modalities (RGB vs grayscale), very long-term time spans, etc., which makes it quite difficult for VPR.
SVOX [11] is a cross-domain VPR dataset collected in a variety of weather and lighting conditions. It includes a large-scale database sourced from Google Street View images spanning the city of Oxford. The queries are extracted from the Oxford RobotCar dataset [44] and divided into multiple subsets for different weather and lighting conditions. We evaluate the model performance using the three most challenging query subsets: SVOX-Night, SVOX-Rain, and SVOX-Sun.
15 Compared Methods Details
NetVLAD [5] is a well-known VPR approach with a differentiable VLAD layer, which can be integrated into common neural networks. In our experiments, we use its PyTorch implementation [^1] https://github.com/Nanne/pytorch-NetVlad with the released VGG16 model trained on Pitts30k for comparison.
SFRS [23] utilizes self-supervised image-to-region similarities to mine hard positive samples for training a more robust NetVLAD model. In the comparison experiments, we follow its official implementation [^2] https://github.com/yxgeee/OpenIBL with the model trained on the Pitts30k dataset.
Patch-NetVLAD [26] is a two-stage method that utilizes NetVLAD-based multi-scale patch-level features to re-rank the candidate images retrieved using NetVLAD global features. The official implementation [^3] https://github.com/QVPR/Patch-NetVLAD with the performance-focused configuration is used in our experiments. Following the original paper, the model trained on the Pitts30k dataset is tested on Pitts30k and Tokyo24/7, while the model trained on the MSLS dataset is evaluated on MSLS (-val and -challenge).
TransVPR [61] is a two-stage VPR method that leverages attentions from three levels of Transformer to produce global features for candidates retrieval, and employs an attention mask to filter feature maps to yield key-patch descriptors for re-ranking candidates. The official implementation [^4] https://github.com/RuotongWANG/TransVPR-model-implementation is used for comparison experiments. The model trained on the Pitts30k dataset is evaluated on Pitts30k and Tokyo24/7, and the model trained on the MSLS dataset is assessed on MSLS.
CosPlace [8] treats VPR model training as a classification problem and trains the model on the individually constructed San Francisco eXtra Large (SF-XL) datasets with the Large Margin Cosine Loss (i.e., cosFace) to achieve remarkable results. We follow its official implementation [^5] https://github.com/gmberton/CosPlace with the VGG16 backbone (producing 512-dim global features) for testing.
GCL [38] uses an automatic annotation strategy producing graded similarity labels for image pairs to re-label VPR datasets, and a novel generalized contrastive loss to utilize such labels to train contrastive networks. we use the results (yield by the version using ResNet152-GeM with PCA) from the original paper for comparison.
MixVPR [2] introduces a novel holistic feature aggregation approach for global-retrieval-based VPR. It utilizes feature maps yielded by a pre-trained backbone as initial feature representations, and employs a sequence of Feature-Mixer modules to incorporate global relationships into each feature map to produce final global features. We follow the official implementation [^6] https://github.com/amaralibey/MixVPR and its best configuration, i.e., using the ResNet50 backbone producing 4096-dim global features, for comparison experiments.
EigenPlaces [10] can be seen as an improvement work on CosPlace. This work trains the networks on images from different viewpoints (of the same place), thus improving the viewpoint robustness of learned global representations. It is the most recent work and achieves the best performance on most VPR datasets. We follow its official implementation [^7] https://github.com/gmberton/EigenPlaces and the configuration using ResNet50 as the backbone to yield 2048-dim features.
Besides, the results on the three challenging datasets (Nordland, AmsterTime, and SVOX) are directly referenced from the EigenPlaces paper [10]. These results are basically consistent with what we reproduced.
16 Additional Qualitative Results and Failure Cases
In the main paper, we have presented a small number of qualitative results to show the robustness of our approach in challenging scenarios. In this section, we add more examples to vividly demonstrate the performance of VPR methods. Fig. 8, Fig. 9, and Fig. 10 show examples on Pitts30k, Tokyo24/7, and MSLS-val, respectively. These examples demonstrate that our method is more robust against variations in conditions and viewpoints, as well as perceptual aliasing, than previous methods. Fig. 11 and Fig. 12 show examples on AmsterTime and Nordland, which demonstrate that our method can correctly recognize place images over long time spans and under extreme environments in general. However, our method also produces erroneous results in a few cases (the last examples of these two figures) when images from different places are very similar, especially when lacking discriminative landmarks. This is a focus to improve our method in the future.
Figure 8: Qualitative results on Pitts30k. The proposed CricaVPR returns the correct database images, while other methods produce wrong results. In these examples, most of the other methods suffer from perceptual aliasing. In the first two examples, all other methods return highly similar but wrong places. In the third example, the buildings on the right of the images returned by TransVPR, CosPlace, and EigenPlaces are highly similar to the building on the right of the query image, indicating that the appearance of this building is not distinguishable enough, making these methods suffer from perceptual aliasing.
Figure 9: Qualitative results on Tokyo24/7. The proposed CricaVPR returns the correct database images, while other methods produce wrong results. In these examples, the main challenges are the variations in lighting conditions across day and night, as well as perceptual aliasing. In the first example, as the query image is a nighttime image, all methods except for ours and EigenPlaces return nighttime but incorrect images. EigenPlaces returns a similar but incorrect image.
Figure 10: Qualitative results on MSLS-val. The proposed CricaVPR returns the correct database images, while other methods produce wrong results. In the first example, NetVLAD returns a highly similar but wrong image. In the second example, the building on the left of the query image is occluded by trees, causing NetVLAD, TransVPR, and CosPlace to return incorrect results with obvious trees on the left side. In the third example, SFRS, CosPlace, and EigenPlaces return database images that are geographically close to the query image but exceed the set threshold (i.e. still wrong).
Figure 11: Qualitative results on AmsterTime. There is a very long time span between the query (grayscale) image and the reference (RGB) image in this dataset. In the first two examples, the proposed CricaVPR returns the right database images, while other methods produce wrong results. In the first example, the discriminative buildings only occupy a small region of the reference image. In the second example, a new building appears in the reference image, and the original building has undergone some modifications. These cause other methods to return incorrect results. In the last example, there are images from different places in the database that are highly similar to the query image, causing none of the methods to retrieve the correct result.
Figure 12: Qualitative results on Nordland. These examples show drastic variations in conditions (season, weather, and lighting). Meanwhile, there are almost no discriminative buildings in the images. These challenges are difficult to address for previous VPR methods, resulting in incorrect results being returned by all of them. Our method gets the right result in the first two examples but fails in the last one.